### v1

@@ Appendix B Adam-based LaPerm algorithm
Here we briefly introduce the Adam-based LaPerm algorithm with a fixed permutation period $k$ in [34]. The pseudocode is shown in Algorithm 1.
@@ Appendix C The impact of permutation period on convergence behavior
As a hyperparameter, the choice of permutation period $k$ during the implementation of LaPerm algorithms has the possibility to affect the convergence behavior. The correlation between the value of $k$ and the final accuracy is reported to be unambiguous (refer to Fig. 6 in [34]). Generally, a larger $k$ is associated with slightly higher accuracy of single permutation training result, thus in our experiments, the weights are permuted after each $k$ epoch. Fig. 7 evaluates the impact of $k$’s value on convergence behavior, whose results suggest that this effect remains negligible.
@@ Appendix E Extending to the networks with deeper architecture or sparse initializations
Additionally, this paper’s conclusions could be extended to sparse initialization scenarios, an area where permutation training shows significant advantages (see Fig. 6(e) in [34]). On an intuitive level, our proof can explain these advantages, as a primary challenge of permutation training is managing unused basis functions. However, in sparse situations, we can conveniently deal with this problem by assigning the remaining basis functions zero coefficients, thereby facilitating the proof process.

### v2

@@ 6 Conclusion
CRediT authorship contribution statement
Y. Cai: Conceptualization, Validation, Formal analysis, Writing - original draft, Writing - review & editing, Visualization, Project administration. G. Chen: Conceptualization, Methodology, Software, Validation, Formal analysis, Investigation, Data curation, Writing - original draft, Writing - review & editing, Visualization. Z. Qiao: Validation, Resources, Writing - review & editing, Supervision, Project administration, Funding acquisition.
Declaration of competing interest
The authors declare that they have no known competing financial interests or personal relationships that could have appeared to influence the work reported in this paper.
Data availability
The code and data accompanying this manuscript are publicly available on GitHub at https://github.com/DanclaChen/PermutationTraining.
Acknowledgments
This work is supported by the CAS AMSS-PolyU Joint Laboratory of Applied Mathematics. The work of Y. Cai is supported by the National Natural Science Foundation of China under Grant 12201053 and 12171043. The work of Z. Qiao is supported by the Hong Kong Research Grants Council RFS grant RFS2021-5S03 and GRF grant 15302122, and the Hong Kong Polytechnic University grant 4-ZZLS.
@@ Appendix B Adam-based LaPerm algorithm
Here we briefly introduce the Adam-based LaPerm algorithm with a fixed permutation period $k$ in Qiu and Suda (2020). The pseudocode is shown in Algorithm 1.
@@ Appendix C The impact of permutation period on convergence behavior
As a hyperparameter, the choice of permutation period $k$ during the implementation of LaPerm algorithms has the possibility to affect the convergence behavior. The correlation between the value of $k$ and the final accuracy is reported to be unambiguous (refer to Fig. 6 in Qiu and Suda (2020)). Generally, a larger $k$ is associated with slightly higher accuracy of single permutation training result, thus in our experiments, the weights are permuted after each $k$ epoch. Fig. 7 evaluates the impact of $k$’s value on convergence behavior, whose results suggest that this effect remains negligible.
@@ Appendix E Extending to the networks with deeper architecture or sparse initializations
Additionally, this paper’s conclusions could be extended to sparse initialization scenarios, an area where permutation training shows significant advantages (see Fig. 6(e) in Qiu and Suda (2020)). On an intuitive level, our proof can explain these advantages, as a primary challenge of permutation training is managing unused basis functions. However, in sparse situations, we can conveniently deal with this problem by assigning the remaining basis functions zero coefficients, thereby facilitating the proof process.
Appendix F A toy example of our step-matching and constant-matching construction
To illustrate the key idea of our construction, we consider a toy example of approaching a step function $f_{s}(x)=0.8\,\chi(x-0.4)$ for $x\in[0,1]$ by a network in Eq. (2) with $n=11$ and equidistantly initialized $\{b_{k}\}_{k=1}^{11}=\{0,0.1,\cdots,1\}$. This setting is considered after applying Lemma 2.1 with $\Delta h=0.8$.
Following our step-matching construction, we can choose the basis functions with $\{b_{k}\}_{k=1}^{4}=\{0.1,0.3,0.6,0.8\}$ to be step-matching in Eq. (6), leading to a step function approximator $f_{s}^{\text{NN}}$. The height of $f_{s}^{\text{NN}}$ is given by $h=0.4$. Therefore, the target step function $f_{s}$ can be approximated by $f_{s}^{\text{NN}}$ along with shifting $\Delta h/2=0.4$ and scaling $\Delta h/h=2$ with the following error estimation:
$\left|\left[\frac{\Delta h}{h}(x)\,f_{s}^{\text{NN}}+\frac{h}{2}\right]-f_{s}(x)\right|\leq 0.8=\Delta h,\quad x\in[0,1].$
The target step function $f_{s}$ and its approximator $f_{s}^{\text{NN}}$ are plotted in Fig. 9.
Figure 9: Approximating a step function $f_{s}(x)=0.8\,\chi(x-0.4)$ and a constant function $f_{c}(x)=0.2$ in $x\in[0,1]$ with the step-function approximator $f_{s}^{\text{NN}}$ and the constant function approximator $f_{c}^{\text{NN}}$, respectively. The target functions are plotted as lines, and the approximation results are shown as dashed lines. The locations of the basis functions used for $f_{s}^{\text{NN}}$ and $f_{c}^{\text{NN}}$ are marked with ”$\times$” and $\blacktriangle$” on the $x$-axis, respectively.
Additionally, we can consider the constant function approximator $f_{c}^{\text{NN}}$ following our constant-matching construction in Eq. (10). We choose the basis functions located at $\{0.2,0.4,0.7,0.9\}$ to be constant-matching, leading to a constant function approximator $f_{c}^{\text{NN}}=h/2$. The height of $f_{c}^{\text{NN}}$ is given by $h/2=0.4$. From the shape of $f_{c}^{\text{NN}}$ plotted in Fig. 9, we can see that it can approximate a constant function without any error.
Appendix G A toy example of Lemma 3.1
To illustrate the processing method of Lemma 3.1, we consider a given sequence
3, 1, 4,1, 5, 9,3141593,\,1,\,4,\,-1,\,5,\,9,3 , 1 , 4 , - 1 , 5 , 9 ,
with the largest gap $\Delta c=4$. The aim is to choose choose $m_{i}\in\{-1,1\}$ to reduce the absolute value of the result
$S=3\cdot m_{1}+1\cdot m_{2}+4\cdot m_{3}-1\cdot m_{4}+5\cdot m_{5}+9\cdot m_{6}.$
This can be achieved by the following steps:
  • 1.
    Rearrange the sequence into 9, 5, 4, 3, 1,19543119,\,5,\,4,\,3,\,1,\,-19 , 5 , 4 , 3 , 1 , - 1;
  • 2.
    Compute the gap between every two adjacent elements as 4, 1, 24124,\,1,\,24 , 1 , 2;
  • 3.
    Reorganize the gap sequence into an alternating series and compute the result as $S=4-2+1=3$;
  • 4.
    Transfer the choice of $m_{i}\in\{-1,1\}$ to the original sequence as
    $S=(9-5)-[1-(-1)]+(4-3)=9-5-1-1+4-3=3<4=\Delta c.$
Therefore, the result $S=3$ can be achieved by choosing $m_{1}=-1,m_{2}=-1,m_{3}=1,m_{4}=1,m_{5}=-1,m_{6}=1$.